# Alberta Wells Dataset: Pinpointing Oil and Gas Wells from Satellite Imagery

### v2

@@ B.1 Dataset Structure
This approach ensures that the training, validation, and test sets include representations from all geographic regions, providing a diverse and comprehensive evaluation. Thus, the dataset represents various geographical regions and offers a diverse benchmark for evaluation and testing.
Each dataset split is saved in an HDF5 format file, structured as described in the following sections, and then compressed into a .tar.gz file for faster transfer. Details on the number of samples in each set and the size of the dataset, both original and compressed, are presented in Table 3.
@@ B.3 Structure of Dataset Directory
To enhance the efficiency of the data loader, we split the larger .h5 dataset into smaller .h5 files, each corresponding to a unique sample (image patch). By splitting the dataset in such a manner, we are able to improve the speed per iteration of the dataloader by over 100%. This results in the following data structure:
@@
B.4 Dataset Size & Distribution of Samples
Our dataset comprises over 94,000 patches of satellite imagery containing wells, with a total of 188,000 patches sourced from Planet Labs [9], covering more than 213,000 individual wells. Details about the distribution of the number of patches, wells present, and dataset split sizes are provided in Table 9, with the distribution of the number of wells per sample being described in Table 10. We also include an equal number of images that contain no wells in each dataset split. The distribution of wells per sample, along with the corresponding number of wells and the breakdown of well types, is illustrated in Figure LABEL:fig:figure_dist_act_sus_abd and detailed in Tables 11, 12, and 13 . The geographic distribution of wells in the dataset can be visualized in Figure 10.
@@ B.5 PlanetScope Satellite Imagery
For our experiments, we selected a 4-band (RGBN) satellite imagery product (ortho_analytic_4b_sr) from Planet Labs [9] as illustrated in Figure LABEL:fig:figure1a. This product uses Planet’s PSB.SD instrument, which features a telescope with a larger 47-megapixel sensor and is designed to be interoperable with Sentinel-2 imagery in several bands. The frequency of each band of image is described in Table 14. The instrument provides a frame size of 32.5 km x 19.6 km, an image capture capacity of 200 million km²/day, and an imagery bit depth of 12-bit, with a ground sample distance (nadir) ranging from 3.7 m to 4.2 m.
@@
Figure 10: Density map of wells in the Alberta Wells Dataset.
@@ Appendix C Dataset Samples Illustration
Samples from the dataset, covering various scenarios, are shown in Figures 13 and 14.
Figure 13: Qualitative results from the dataset illustrate the diverse distribution of wells in dataset samples, including Bbox annotations and corresponding imagery in different spectral bands.
Figure 14: The qualitative results from the dataset showcase the varied distribution of wells in dataset samples, with their corresponding segmentation labels (binary and multi-class) and Bbox annotations.
@@ E.1 Motivation
    The raw data is sourced from the Alberta Energy Regulator (AER), specifically from the monthly AER ST37 publication. This dataset includes comprehensive details about all reported wells in Alberta, such as their geographic location, mode of operation, license status, and the type of product extracted, among other attributes. The data is provided in shapefile format along with accompanying metadata. However, the dataset cannot be used directly because the license status or mode of operation often does not reflect the well’s actual status. Therefore, the authors include domain experts from Anonymous, who specialize in field measurements of methane and air pollutant emissions from oil, gas, and urban systems, as well as in the geospatial and statistical data analysis of emissions and energy infrastructure, to ensure the quality of the dataset.
@@
    This project was funded by Anonymous.
@@ E.2 Composition
    The satellite imagery used in this project was obtained under Planet Labs’ [9] Education & Research license, which prohibits sharing raw satellite imagery. We re-projected the raw data to EPSG:32611 using the nearest resampling method and removed all geographic metadata, such as image bounds and coordinates, from the shared data imagery to create a derived product that complies with the license terms.
@@ E.3 Collection Process
    We acquired the data from the Alberta Energy Regulator, specifically from its monthly well bulletin AER ST37 [31], dated March 2024. Leveraging domain expertise, we filtered the data to reflect the condition of wells as of September 30, 2023. This decision was made because imagery acquired from Alberta during the winter months tends to have high cloud cover. Therefore, we filtered the data to ensure we could collect the best data for each patch based on satellite data acquired between the summer months of June and September in the region.
@@ E.5 Uses
    This dataset focuses on Alberta, Canada, known for its diverse oil reserves and varied landscapes, providing a representative sample comparable to regions in the Appalachian and Mountain West areas of the United States and some former Soviet states with oil wells and unidentified site issues. A limitation of our study is the reliance on well locations from the Alberta Energy Regulator, which may miss some sites, leading to potential false negatives in the ground-truth data. However, this should have minimal impact on algorithm training, as these labels are a minor part of the dataset, and deep learning algorithms can handle moderate label noise well (see e.g., [48]). The main effect may be underreported test accuracy, with some correctly predicted well locations wrongly counted as false. We plan to investigate this further in future work. Additionally, the use of multi-spectral optical data in the AWD dataset may limit the models’ applicability in regions with frequent cloud cover.
@@ E.6 Distribution
    The Dataset is made public and hosted on Zenodo at https://github.com/RolnickLab/Alberta_Wells_Dataset .
@@
    The satellite imagery used in this project was acquired under the Education & Research license of Planet Labs [9]. This license allows for the use of the data in publications and the creation of derivative products, which can be shared in association with publications. However, raw imagery cannot be shared publicly. To comply with these guidelines, we share the data in HDF5 format, with satellite imagery represented as a numpy array from Raster Vector. We have removed all geographic metadata, such as image bounds and coordinates, from the shared data. The data is intended for academic use only and should not be used for commercial purposes. Proper credit must be given to the current authors, Planet Labs, and the Alberta Energy Regulator when using this data.

### v3

@@ B.1 Dataset Structure
This approach ensures that the training, validation, and test sets include representations from all geographic regions, providing a diverse and comprehensive evaluation. Thus, the dataset represents various geographical regions and offers a diverse benchmark for evaluation and testing. Each dataset split is saved in an HDF5 format file, structured as described in the following sections, and then compressed into a .tar.gz file for faster transfer.
Details on the number of samples in each set and the size of the dataset, both original and compressed, are presented in Table 3.
@@ B.3 Structure of Dataset Directory
To enhance the efficiency of the data loader, we split the larger .h5 dataset into smaller .h5 files, each corresponding to a unique sample (image patch). By splitting the dataset in such a manner, we are able to improve the speed per iteration of the dataloader by over 100%.
This results in the following data structure:
@@
B.4 Dataset Size & Distribution of Samples
Our dataset comprises over 94,000 patches of satellite imagery containing wells, with a total of 188,000 patches sourced from Planet Labs (PBC, 2024), covering more than 213,000 individual wells. Details about the distribution of the number of patches, wells present, and dataset split sizes are provided in Table 9, with the distribution of the number of wells per sample being described in Table 10. We also include an equal number of images that contain no wells in each dataset split. The distribution of wells per sample, along with the corresponding number of wells and the breakdown of well types, is illustrated in Figure LABEL:fig:figure_dist_act_sus_abd and detailed in Tables 11, 12, and 13 . The geographic distribution of wells in the dataset can be visualized in Figure 10.
@@ B.4 Dataset Size & Distribution of Samples
Figure 10: Density map of wells in the Alberta Wells Dataset.
@@ B.5 PlanetScope Satellite Imagery
For our experiments, we selected a 4-band (RGBN) satellite imagery product (ortho_analytic_4b_sr) from Planet Labs (PBC, 2024) as illustrated in Figure LABEL:fig:figure1a. This product uses Planet’s PSB.SD instrument, which features a telescope with a larger 47-megapixel sensor and is designed to be interoperable with Sentinel-2 imagery in several bands. The frequency of each band of image is described in Table 14. The instrument provides a frame size of 32.5 km x 19.6 km, an image capture capacity of 200 million km²/day, and an imagery bit depth of 12-bit, with a ground sample distance (nadir) ranging from 3.7 m to 4.2 m.
@@ Appendix C Dataset Samples Illustration
Samples from the dataset, covering various scenarios, are shown in Figures 16 and 17. A few more samples with prediction of the models and labels for binary segmentation and binary detection is illustrated in Figure 18.
Figure 16: Qualitative results from the dataset illustrate the diverse distribution of wells in dataset samples, including
Bbox annotations and corresponding imagery in different spectral bands.
Figure 17: The qualitative results from the dataset showcase the varied distribution of wells in dataset samples, with their
corresponding segmentation labels (binary and multi-class) and Bbox annotations.
@@ Appendix D Challenges for ML Community
Figure 18: Few sample image patches from our dataset includes examples with no wells, two wells, and multiple wells. Additionally, we present qualitative results with predictions generated by our Segmentation U-Net (ResNet50) and DETR ResNet50 model.
@@ E.1 Motivation
    The raw data is sourced from the Alberta Energy Regulator (AER), specifically from the monthly AER ST37 publication. This dataset includes comprehensive details about all reported wells in Alberta, such as their geographic location, mode of operation, license status, and the type of product extracted, among other attributes. The data is provided in shapefile format along with accompanying metadata. However, the dataset cannot be used directly because the license status or mode of operation often does not reflect the well’s actual status. Therefore, the authors include domain experts, who specialize in field measurements of methane and air pollutant emissions from oil, gas, and urban systems, as well as in the geospatial and statistical data analysis of emissions and energy infrastructure, to ensure the quality of the dataset.
@@
    This work was funded in part through the Canada CIFAR AI Chairs program and a McGill Tomlinson Scientist Award. We acknowledge support from Planet Labs in the form of their academic subscription program, and are grateful for the help of Tim Elrick in working with Planet data. We also thank Mila and the Digital Research Alliance of Canada for provision of computing facilities, and NVIDIA Corporation for additional computational resources.
@@ E.2 Composition
    The satellite imagery used in this project was obtained under Planet Labs’ (PBC, 2024) Education & Research license, which prohibits sharing raw satellite imagery. We re-projected the raw data to EPSG:32611 using the nearest resampling method and removed all geographic metadata, such as image bounds and coordinates, from the shared data imagery to create a derived product that complies with the license terms.
@@ E.3 Collection Process
    We acquired the data from the Alberta Energy Regulator, specifically from its monthly well bulletin AER ST37 (AER, 2024), dated March 2024. Leveraging domain expertise, we filtered the data to reflect the condition of wells as of September 30, 2023. This decision was made because imagery acquired from Alberta during the winter months tends to have high cloud cover. Therefore, we filtered the data to ensure we could collect the best data for each patch based on satellite data acquired between the summer months of June and September in the region.
@@ E.5 Uses
    This dataset focuses on Alberta, Canada, known for its diverse oil reserves and varied landscapes, providing a representative sample comparable to regions in the Appalachian and Mountain West areas of the United States and some former Soviet states with oil wells and unidentified site issues. A limitation of our study is the reliance on well locations from the Alberta Energy Regulator, which may miss some sites, leading to potential false negatives in the ground-truth data. However, this should have minimal impact on algorithm training, as these labels are a minor part of the dataset, and deep learning algorithms can handle moderate label noise well (see e.g., (Rolnick et al., 2017)). The main effect may be underreported test accuracy, with some correctly predicted well locations wrongly counted as false. We plan to investigate this further in future work. Additionally, the use of multi-spectral optical data in the AWD dataset may limit the models’ applicability in regions with frequent cloud cover.
@@ E.6 Distribution
    The Dataset is made public and hosted on Zenodo at https://zenodo.org/records/13743323 .
@@
    The satellite imagery used in this project was acquired under the Education & Research license of Planet Labs (PBC, 2024). This license allows for the use of the data in publications and the creation of derivative products, which can be shared in association with publications. However, raw imagery cannot be shared publicly. To comply with these guidelines, we share the data in HDF5 format, with satellite imagery represented as a numpy array from Raster Vector. We have removed all geographic metadata, such as image bounds and coordinates, from the shared data. The data is intended for academic use only and should not be used for commercial purposes. Proper credit must be given to the current authors, Planet Labs, and the Alberta Energy Regulator when using this data.